OpenStudy (anonymous):

A circle is defined by the equation x2 + y2 – x – 2y – 11/4 = 0. What are the coordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

\[ x^2 + y^2 – x – 2y – 11/4 =x^2 -x +\frac 1 4 -\frac 14 + y^2 - 2y + 4-4 -11/4=\\ (x-\frac 1 2)^2+ (y-2)^2-\frac 1 4 -4 -\frac{11}4=\\ (x-\frac 1 2)^2+ (y-2)^2-7=0\\ (x-\frac 1 2)^2+ (y-2)^2 =(\sqrt 7)^2 \] Center(1/2 ,2) radius = \sqrt 7

OpenStudy (anonymous):

radius=\( \sqrt 7\)

OpenStudy (anonymous):

(1,1), 2 units?

OpenStudy (anonymous):

Center is \( \left( \frac 1 2 ,2 \right)\)

OpenStudy (anonymous):

radius=\( \sqrt 7\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

YW